# VisJudge-Bench: Aesthetics and Quality Assessment of Visualizations

### v2

@@ 1 Introduction
Visualization serves as an effective approach for transforming complex datasets into intuitive insights (Shen et al., 2023; Qin et al., 2020; Ye et al., 2024; Li et al., 2024a; 2025a). The value of a high-quality visualization depends on whether its data is faithfully presented (Fidelity), whether information is clearly communicated (Expressiveness), and whether the design is aesthetically well-presented (Aesthetics), as shown in Figure 1. These three dimensions are closely interconnected and indispensable, posing challenges for visualization quality assessment.
Although Multimodal Large Language Models (MLLMs) have shown potential in aesthetic evaluation of natural images (Murray et al., 2012; Li et al., 2024c), applying them to visualization evaluation faces unique challenges. Unlike natural images, visualization evaluation requires simultaneous judgment of data encoding accuracy, information communication effectiveness, and visual design appropriateness, as shown in Figure 2. However, existing MLLMs benchmarks are insufficient for such comprehensive evaluation, as detailed in Table 1. First, chart question answering benchmarks (e.g., ChartInsights (Wu et al., 2024b)) evaluate models’ ability to understand chart information, rather than their overall design quality. Second, natural image aesthetic evaluation benchmarks (e.g., ArtiMuse (Li et al., 2024c)) focus on assessing aesthetics, but ignore the core purpose of visualization to effectively communicate data. Finally, existing visualization evaluation benchmarks (e.g., VisEval (Fu et al., 2024)) mainly evaluate natural language to visualization (NL2VIS) tasks (Luo et al., 2021c), with the focus on assessing whether generated visualizations accurately reflect natural language queries, rather than the aesthetics and quality of visualizations. This leads to a critical research gap: we lack a systematic framework to measure MLLMs’ comprehensive capabilities in evaluating visualization aesthetics and quality.
@@
To address this challenge, we construct VisJudge-Bench, the first comprehensive benchmark based on the “Fidelity, Expressiveness, and Aesthetics” principles to assess MLLMs’ capabilities in visualization aesthetics and quality evaluation. It contains 3,090 expert-scored samples from real-world scenarios, covering single visualizations, multiple visualizations, and dashboards across 32 chart types. Using this benchmark, we conduct extensive testing on multiple MLLMs, including GPT-5, finding that even the most advanced models show significant differences from human experts (Mean Absolute Error as high as 0.553, correlation only 0.428). This finding clearly demonstrates that general MLLMs cannot automatically acquire specialized evaluation capabilities in the visualization domain, making the development of specialized optimization models necessary.
@@ 2 Related Work
MLLM as a Judge. Recently, MLLMs have shown significant potential in emulating human expert judgment, a paradigm known as “MLLM-as-a-Judge” (Zheng et al., 2023; Chen et al., 2024). As summarized in Table 1, these works can be categorized into three groups. The first is general visual aesthetics assessment, where models evaluate the artistic quality of photographs, as seen in AVA (Murray et al., 2012) and ArtiMuse (Cao et al., 2025). However, this overlooks the critical aspect of information communication efficiency in data visualization. The second is chart understanding tasks, with examples like ChartQA (Masry et al., 2022) and ChartInsights (Wu et al., 2024b), which assess the model’s ability to understand and interpret chart information. Yet, these works only focus on the ability to “read” chart information. The third is visualization evaluation, where recent works like VisEval (Fu et al., 2024) and VIS-Shepherd (Pan et al., 2025) have explored using MLLMs to judge visualizations in the context of NL2VIS tasks, focusing on whether the chart accurately reflects the natural language query. However, they fall short of a comprehensive evaluation of the intrinsic “design quality”. This reveals a gap in existing research: the absence of a multi-dimensional framework for evaluating data fidelity, information effectiveness, and visual aesthetics. To address this, we introduce VisJudge-Bench, the first comprehensive benchmark designed to systematically evaluate the capabilities of MLLMs as “visualization quality judges”.
@@
ArtiMuse (Li et al., 2024c) Images General Images × ×
@@
ChartInsights (Wu et al., 2024b) Chart, Question Single Vis × ×
Visualization Evaluation VisEval (Fu et al., 2024) Chart, NL, Data Single Vis × ×
@@ 3.1.1 Data Collection and Preprocessing
This stage constructs the visualization corpus through two key components: corpus construction and data preprocessing.
@@ 3.1.2 The “Fidelity, Expressiveness, and Aesthetics” Evaluation Framework
The “Fidelity, Expressiveness, and Aesthetics” Framework Design. To systematically evaluate visualization quality, we construct a multi-dimensional evaluation framework. This framework draws inspiration from classical translation theory principles of “Fidelity, Expressiveness, and Aesthetics” (illustrated with positive-negative examples in Figure 1), combined with established theories in graphical perception (Cleveland and McGill, 1984), information visualization design (Munzner, 2014), and aesthetic evaluation (Li et al., 2024c). We operationalize this core concept into six measurable evaluation dimensions (as shown in Figure 3):
@@
  • Fidelity focuses on Data Fidelity. This dimension draws from Tufte’s design principles (Tufte, 1983) for avoiding “graphical lies” and recent research on visualization misleadingness issues (Nguyen et al., 2013; Szafir, 2018; Lan and Liu, 2024b; McNutt et al., 2020). Since source data for web-collected visualizations is typically unavailable, this dimension evaluates data presentation accuracy at the visual level. It assesses whether visual encodings accurately reflect the displayed values, examining visually detectable distortions such as improper axis settings, scale distortions, truncated baselines, disproportional encodings, or other misleading design patterns.
@@
  • Expressiveness focuses on the effectiveness of information communication. This dimension evaluates how effectively visualizations convey information to users. It includes two progressive sub-dimensions: First, (1) Semantic Readability evaluates the clarity of basic information encoding, assessing whether users can unambiguously decode visual elements in charts (Pan et al., 2025; Tang et al., 2026). Building on chart readability, (2) Insight Discovery further evaluates the analytical value in revealing deep data patterns, trends, or outliers, helping users transition from “reading information” to “gaining insights” (Wu et al., 2024a; Chen et al., 2025).
@@
  • Aesthetics focuses on Aesthetic Quality of visual design, integrating visualization perception theory (Ware, 2021) with design practice. This dimension consists of three sub-dimensions that collectively influence the overall visual experience: (1) Design Style evaluates the innovation and uniqueness of design, measuring the degree of novel visual elements and distinctive style (Dibia, 2023; Brath and Banissi, 2016); (2) Visual Composition focuses on the rationality of spatial layout, evaluating the balance and order of element positioning, size proportions, and spacing arrangements (Wu et al., 2023); and (3) Color Harmony evaluates the coordination and functionality of color combinations, ensuring color palette choices balance aesthetics with effective information communication (Harrower and Brewer, 2003; Gramazio et al., 2017).
@@
In addition, this evaluation framework offers flexibility, with specific evaluation criteria and score weights adaptively customized according to different visualization types (such as single visualizations (Wu et al., 2024a), multiple visualizations (Chen et al., 2020), and dashboards (Bach et al., 2023)). Complete evaluation rules and customization details are provided in Appendix C.
@@
Adaptive Question Generation Mechanism. Based on the evaluation framework, we have devised an adaptive question generation process (detailed workflow shown in Figure 3). This process begins by leveraging GPT-4o to extract metadata from the chart, such as its type and visual elements. Subsequently, it rewrites questions by populating predefined templates based on this metadata, generating highly customized questions and their corresponding five-point scoring criteria for the six evaluation sub-dimensions. For instance, under the data fidelity dimension, the generated question specifically targets the chart’s data presentation accuracy, asking whether the “horizontal bar chart accurately represents injury counts with a proper zero baseline and proportional bar lengths.” The scoring criteria focuses on visually detectable issues: a score of 1 identifies “major distortions, such as truncated axes or misleading scales that exaggerate differences,” whereas a score of 5 confirms a “highly faithful representation where bar lengths are strictly proportional to the displayed values.” This approach ensures that the evaluation questions are closely aligned with the specific visualization content. For more detailed examples, please refer to Appendix C.1.
@@ 3.1.3 Expert Annotation and Quality Control
To build reliable human ground truth, VisJudge-Bench adopts a rigorous three-stage annotation and quality control workflow (bottom panel of Figure 3) informed by benchmark construction approaches (Rein et al., 2024; Liu et al., 2025; Zhu et al., 2024). This systematic process ensures high-fidelity and consistent scoring through careful review and expert judgment.
Stage 1: Initial Annotation. We recruited 603 highly qualified crowdsourcing workers through the CloudResearch platform (CloudResearch, 2022). To ensure annotation quality, we set strict screening criteria (see Appendix A.3.1 for details) and designed a dedicated annotation interface (Appendix A.3.3). Crucially, we embedded validation checks to identify and filter out inattentive responses (examples in Appendix A.3.4). Each of the 3,090 samples was scored by three independent annotators across six evaluation dimensions (task design details in Appendix A.3.2), generating an initial scoring matrix.
Stage 2: Quality Control. To address scoring disagreements among annotators, we designed a systematic conflict identification and resolution mechanism based on established crowdsourcing quality control and statistical evaluation theory (Gadiraju et al., 2015; Rousseeuw and Leroy, 2005; Brennan, 2001). The system first identifies high-disagreement samples by analyzing score variance, then algorithmically generates candidate resolution strategies including outlier removal, malicious scoring detection, and sub-dimensional bias correction. These algorithm-generated suggestions are processed and ranked before being submitted to the expert team for final review (complete algorithmic details and parameters in Appendix A.3.4).
@@ 4 VisJudge: A Specialized Model for Visualization Evaluation
Model Training. We fine-tuned four representative open-source multimodal models as base models: Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-3B-Instruct (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025), and Llava-v1.6-mistral-7B (Liu et al., 2024). These models span different architectures and parameter scales (3B to 8B), enabling comprehensive evaluation of our training approach’s cross-architecture generalization. All models are trained to generate quality scores (1.0-5.0) and rationales aligned with human expert judgments. We employ reinforcement learning with the GRPO algorithm (Shao et al., 2024), using a composite reward function combining accuracy reward (minimizing prediction error) and format reward (ensuring structured outputs) (Shi et al., 2025; Wu et al., 2025). Formal reward definitions are detailed in Appendix D.2. For parameter-efficient fine-tuning, we adopted Low-Rank Adaptation (LoRA) (Hu et al., 2022). All models were trained using consistent hyperparameter configurations across architectures. Training used 5 epochs with a learning rate of 1e-5. Detailed configurations are in Appendix D.3.
@@ 5.1 Experimental Settings
Evaluation Setup. We evaluate 12 representative MLLMs: GPT-5, GPT-4o, Claude-4-Sonnet, Claude-3.5-Sonnet, Gemini-2.0-Flash, Gemini-2.5-Pro, Qwen2.5-VL (3B/7B/32B/72B-Instruct) (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025), Llava-v1.6-mistral-7B (Liu et al., 2024), and their corresponding fine-tuned variants, which we collectively refer to as VisJudge on a balanced test set of 648 samples (see Appendix A.2.7 for distribution details). Each model provides 1-to-5 scores with justifications based on our evaluation framework. Following human annotation procedures, we run each model three times and average the results. All inference uses vLLM on four NVIDIA A6000 (48GB) GPUs with bfloat16 precision and a temperature of 0.8.
@@ 5.2.1 Can MLLMs Assess Visualization Aesthetics and Quality Like Humans?
Hierarchical Capability Structure. Current MLLMs exhibit a clear hierarchical performance structure across evaluation dimensions. Models perform relatively well on “Fidelity” dimensions (average MAE 1.006), reflecting their fundamental capability in identifying obvious data errors (e.g., proportion distortions, baseline issues). Within “Expressiveness” dimensions, models show better performance on Insight Discovery (average MAE 0.883) compared to Semantic Readability (average MAE 0.991). Most prominently, all models struggle significantly with “Aesthetics” dimensions across all three aesthetic sub-dimensions, with average MAE around 0.755 and correlations of 0.408, 0.177, and 0.220 for Design Style, Composition, and Color respectively, highlighting the inherent challenges of subjective aesthetic assessment, which often involves nuanced cultural context and abstract design principles that are difficult for current models to grasp.
Model-Specific Evaluation Characteristics. Through fine-grained analysis, we identify distinct “evaluation personalities” across different models. GPT-5 demonstrates balanced performance across dimensions with consistently competitive scores, particularly excelling in overall accuracy; GPT-4o shows relative strength in Color Harmony assessment (MAE 0.657), reflecting sensitivity to color aesthetics; Claude-4-Sonnet excels in Semantic Readability evaluation (MAE 0.757), showing advantages in information communication assessment; while Gemini-2.0-Flash leads in Data Fidelity (MAE 0.829), indicating focus on data accuracy. Among open-source models, Qwen2.5-VL-72B-Instruct achieves the best overall performance (MAE 0.702, Corr. 0.440), approaching commercial model capabilities, while smaller models like Qwen2.5-VL-3B-Instruct and Llava-v1.6-mistral-7B show larger performance gaps. These differentiated capability distributions validate VisJudge-Bench’s diagnostic value and provide guidance for practical model selection.
Domain-Specific Fine-tuning Effectiveness. Our specialized VisJudge achieves superior performance across all core metrics. The best-performing variant based on Qwen2.5-VL-7B-Instruct reaches an overall MAE of 0.421 and correlation of 0.687. Among commercial models, GPT-5 achieves the best MAE performance (0.553) while GPT-4o reaches the highest correlation (0.482). Compared to these strong baselines, VisJudge demonstrates substantial improvements: 23.9% MAE reduction over GPT-5 (from 0.553 to 0.421) and 42.5% correlation improvement over GPT-4o (from 0.482 to 0.687), demonstrating the substantial potential of domain-specific fine-tuning. Notably, domain-specific fine-tuning consistently reduces MAE across different backbone architectures, yielding 30–40% error reductions on all three open-source families. Qwen2.5-VL-3B-Instruct achieves a 40.2% reduction (from 0.821 to 0.491), InternVL3-8B 31.8% (0.793 to 0.541), and Llava-v1.6-mistral-7B 31.5% (0.724 to 0.496), demonstrating robust generalization of our training methodology across diverse model families. In subsequent analyses, VisJudge refers to this best-performing variant (Qwen2.5-VL-7B-Instruct based) unless otherwise specified.

### v3

@@ 1 Introduction
Visualization serves as an effective approach for transforming complex datasets into intuitive insights (Shen et al., 2023; Qin et al., 2020; Ye et al., 2024; Li et al., 2024a; Zhu et al., 2025b; Li et al., 2025b; a). The value of a high-quality visualization depends on whether its data is faithfully presented (Fidelity), whether information is clearly communicated (Expressiveness), and whether the design is aesthetically well-presented (Aesthetics), as shown in Figure 1. These three dimensions are closely interconnected and indispensable, posing challenges for visualization quality assessment.
Although Multimodal Large Language Models (MLLMs) have shown potential in aesthetic evaluation of natural images (Murray et al., 2012; Cao et al., 2025), applying them to visualization evaluation faces unique challenges. Unlike natural images, visualization evaluation requires simultaneous judgment of data encoding accuracy, information communication effectiveness, and visual design appropriateness, as shown in Figure 2. However, existing MLLM benchmarks are insufficient for such comprehensive evaluation, as detailed in Table 1. First, chart question answering benchmarks (e.g., ChartInsights (Wu et al., 2024)) evaluate models’ ability to understand chart information, rather than their overall design quality. Second, natural image aesthetic evaluation benchmarks (e.g., ArtiMuse (Cao et al., 2025)) focus on assessing aesthetics, but ignore the core purpose of visualization to effectively communicate data. Finally, existing visualization evaluation benchmarks (e.g., VisEval (Chen et al., 2024b)) mainly evaluate natural language to visualization (NL2VIS) tasks (Luo et al., 2021c), with the focus on assessing whether generated visualizations accurately reflect natural language queries, rather than the aesthetics and quality of visualizations. This leads to a critical research gap: we lack a systematic framework to measure MLLMs’ comprehensive capabilities in evaluating visualization aesthetics and quality.
@@
To address this challenge, we construct VisJudge-Bench, the first comprehensive benchmark based on the “Fidelity, Expressiveness, and Aesthetics” principles to assess MLLMs’ capabilities in visualization aesthetics and quality evaluation. It contains 3,090 expert-scored samples from real-world scenarios, covering single visualizations, multiple visualizations, and dashboards across 32 chart types. Using this benchmark, we conduct extensive testing on 12 representative MLLMs, including GPT-5, finding that even the most advanced models show significant differences from human experts (MAE as high as 0.553, correlation only 0.428). This finding clearly demonstrates that general MLLMs cannot automatically acquire specialized evaluation capabilities in the visualization domain, making the development of specialized optimization models necessary.
@@ 2 Related Work
MLLM as a Judge. Recently, MLLMs have shown significant potential in emulating human expert judgment, a paradigm known as “MLLM-as-a-Judge” (Zheng et al., 2023; Chen et al., 2024a; Bian et al., 2025). As summarized in Table 1, these works fall into three categories: (1) general visual aesthetics assessment (e.g., AVA (Murray et al., 2012), ArtiMuse (Cao et al., 2025)), which evaluates the artistic quality of photographs but overlooks information communication efficiency in data visualization; (2) chart understanding tasks (e.g., ChartQA (Masry et al., 2022), ChartInsights (Wu et al., 2024)), which focus on understanding and interpreting chart content but neglect overall design quality; and (3) visualization evaluation (e.g., VisEval (Chen et al., 2024b), VIS-Shepherd (Pan et al., 2025)), which evaluates whether generated visualizations accurately reflect natural language queries, but lacks a comprehensive assessment of intrinsic design quality. To address this gap, we introduce VisJudge-Bench, the first comprehensive benchmark designed to systematically evaluate the capabilities of MLLMs as “visualization quality judges”.
@@
ArtiMuse (Cao et al., 2025) Images General Images × ×
@@
ChartInsights (Wu et al., 2024) Chart, Question Single Vis × ×
Visualization Evaluation VisEval (Chen et al., 2024b) Chart, NL, Data Single Vis × ×
@@ 3.1.2 The “Fidelity, Expressiveness, and Aesthetics” Evaluation Framework
The “Fidelity, Expressiveness, and Aesthetics” Framework Design. To systematically evaluate visualization quality, we construct a multi-dimensional evaluation framework. This framework draws inspiration from classical translation theory principles of “Fidelity, Expressiveness, and Aesthetics” (illustrated with positive-negative examples in Figure 1), combined with established theories in graphical perception (Cleveland and McGill, 1984), information visualization design (Munzner, 2025), and aesthetic evaluation (Cao et al., 2025). We operationalize this core concept into six measurable evaluation dimensions (as shown in Figure 3):
@@
  • Fidelity focuses on Data Fidelity. This dimension draws from Tufte’s design principles (Tufte, 1983) for avoiding “graphical lies” and recent research on visualization misleadingness issues (Nguyen et al., 2013; Szafir, 2018; Lan and Liu, 2024; McNutt et al., 2020; Xie et al., 2026). Since source data for web-collected visualizations is typically unavailable, this dimension evaluates data presentation accuracy at the visual level. It assesses whether visual encodings accurately reflect the displayed values, examining visually detectable distortions such as improper axis settings, scale distortions, truncated baselines, or other misleading design patterns.
@@
  • Expressiveness focuses on the effectiveness of information communication. This dimension evaluates how effectively visualizations convey information to users. It includes two progressive sub-dimensions: First, (1) Semantic Readability evaluates the clarity of basic information encoding, assessing whether users can unambiguously decode visual elements in charts (Pan et al., 2025; Tang et al., 2026; Wang et al., 2026). Building on chart readability, (2) Insight Discovery further evaluates the analytical value in revealing deep data patterns, trends, or outliers, helping users transition from “reading information” to “gaining insights” (Chen et al., 2024b; 2025).
@@
  • Aesthetics focuses on Aesthetic Quality of visual design, integrating visualization perception theory (Ware, 2019) with design practice. This dimension consists of three sub-dimensions that collectively influence the overall visual experience: (1) Design Style evaluates the innovation and uniqueness of design, measuring the degree of novel visual elements and distinctive style (Dibia, 2023; Brath and Banissi, 2016); (2) Visual Composition focuses on the rationality of spatial layout, evaluating the balance and order of element positioning, size proportions, and spacing arrangements (Wu et al., 2023); and (3) Color Harmony evaluates the coordination and functionality of color combinations, ensuring color palette choices balance aesthetics with effective information communication (Harrower and Brewer, 2003; Gramazio et al., 2017).
@@
In addition, this evaluation framework offers flexibility, with evaluation questions and scoring criteria adaptively tailored to each visualization’s type (such as single visualizations (Chen et al., 2024b), multiple visualizations (Chen et al., 2020), and dashboards (Bach et al., 2023)) and its specific visual characteristics. Complete evaluation rules and customization details are provided in Appendix C.
@@
Adaptive Question Generation Mechanism. Based on the evaluation framework, we have devised an adaptive question generation process (detailed workflow shown in Figure 3). This process begins by leveraging GPT-4o to extract metadata from the chart, such as its type and visual elements. Subsequently, it rewrites questions by populating predefined templates based on this metadata, generating highly customized questions and their corresponding five-point scoring criteria for the six evaluation sub-dimensions. For instance, under the data fidelity dimension, the generated question specifically targets the chart’s data presentation accuracy, asking whether the “horizontal bar chart accurately represents injury counts with a proper zero baseline and proportional bar lengths.” The scoring criteria focus on visually detectable issues: a score of 1 identifies “major distortions, such as truncated axes or misleading scales that exaggerate differences,” whereas a score of 5 confirms a “highly faithful representation where bar lengths are strictly proportional to the displayed values.” This approach ensures that the evaluation questions are closely aligned with the specific visualization content. For more detailed examples, please refer to Appendix C.1.
@@ 3.1.3 Expert Annotation and Quality Control
To build reliable human ground truth, VisJudge-Bench adopts a rigorous three-stage annotation and quality control workflow (bottom panel of Figure 3) informed by benchmark construction approaches (Rein et al., 2024; Liu et al., 2025; Zhu et al., 2024; Shen et al., 2026). This systematic process ensures high-fidelity and consistent scoring through careful review and expert judgment.
Stage 1: Initial Annotation. We recruited 603 highly qualified crowdsourcing workers through the CloudResearch platform (Hartman et al., 2023). To ensure annotation quality, we set strict screening criteria (see Appendix A.3.1 for details) and designed a dedicated annotation interface (Appendix A.3.3). Crucially, we embedded validation checks to identify and filter out inattentive responses (examples in Appendix A.3.4). Each of the 3,090 samples was scored by three independent annotators across six evaluation dimensions (task design details in Appendix A.3.2), generating an initial scoring matrix.
Stage 2: Quality Control. To address scoring disagreements among annotators, we designed a systematic conflict identification and resolution mechanism based on established crowdsourcing quality control and statistical evaluation theory (Gadiraju et al., 2015; Rousseeuw and Leroy, 2003; Brennan, 1992). The system first identifies high-disagreement samples by analyzing score variance, then algorithmically generates candidate resolution strategies including outlier removal, malicious scoring detection, and sub-dimensional bias correction. These algorithm-generated suggestions are processed and ranked before being submitted to the expert team for final review (complete algorithmic details and parameters in Appendix A.3.4).
@@ 4 VisJudge: A Specialized Model for Visualization Evaluation
Model Training. We fine-tuned four representative open-source multimodal models as base models: Qwen2.5-VL (3B/7B-Instruct) (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025a), and Llava-v1.6-mistral-7B (Liu et al., 2024). These models span different architectures and parameter scales (3B to 8B), enabling comprehensive evaluation of our training approach’s cross-architecture generalization (Lin et al., 2025). All models are trained to generate quality scores (1.0–5.0) and rationales aligned with human expert judgments. We employ reinforcement learning with the GRPO algorithm (Shao et al., 2024; Liang et al., 2026; Wu et al., 2026), using a composite reward function combining accuracy reward (minimizing prediction error) and format reward (ensuring structured outputs) (Shi et al., 2025; Wu et al., 2025). Formal reward definitions are detailed in Appendix D.2. For parameter-efficient fine-tuning, we adopted Low-Rank Adaptation (LoRA) (Hu et al., 2022). All models were trained using consistent hyperparameter configurations across architectures. Training used 5 epochs with a learning rate of 1e-5. Detailed configurations are in Appendix D.3.
@@ 5.1 Experimental Settings
Evaluation Setup. We evaluate 12 representative MLLMs: GPT-5, GPT-4o, Claude-4-Sonnet, Claude-3.5-Sonnet, Gemini-2.0-Flash, Gemini-2.5-Pro, Qwen2.5-VL (3B/7B/32B/72B-Instruct) (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025a), Llava-v1.6-mistral-7B (Liu et al., 2024), and their corresponding fine-tuned variants, which we collectively refer to as VisJudge, on a balanced test set of 648 samples (see Appendix A.2.7 for distribution details). Each model provides 1-to-5 scores with justifications based on our evaluation framework. Following human annotation procedures, we run each model three times and average the results. All inference uses vLLM on four NVIDIA A6000 (48GB) GPUs with bfloat16 precision and a temperature of 0.8.
@@ 5.2.1 Can MLLMs Assess Visualization Aesthetics and Quality Like Humans?
Hierarchical Capability Structure. Current MLLMs exhibit a hierarchical performance pattern: “Fidelity” is relatively strong, “Expressiveness” is moderate, and “Aesthetics” is the most challenging, with average MAE around 0.755 and correlations of only 0.177–0.408 across the three aesthetic sub-dimensions, reflecting the difficulty of abstract design principles and cultural context for current models. Models also exhibit distinct evaluation characteristics: GPT-5 performs consistently across dimensions; GPT-4o is relatively strong in Color Harmony (MAE 0.657); Claude-4-Sonnet excels in Semantic Readability (MAE 0.757); and Gemini-2.0-Flash leads in Data Fidelity (MAE 0.829). Among open-source models, Qwen2.5-VL-72B-Instruct achieves the best overall performance (MAE 0.702, Corr. 0.440), approaching commercial model capabilities.
Domain-Specific Fine-tuning Effectiveness. VisJudge (based on Qwen2.5-VL-7B-Instruct) achieves the best overall performance with MAE 0.421 and correlation 0.687, representing a 23.9% MAE reduction over GPT-5 and a 42.5% correlation improvement over GPT-4o. Domain-specific fine-tuning consistently yields 30–40% error reductions across different backbone architectures: Qwen2.5-VL-3B-Instruct achieves a 40.2% reduction (from 0.821 to 0.491), InternVL3-8B 31.8% (0.793 to 0.541), and Llava-v1.6-mistral-7B 31.5% (0.724 to 0.496), demonstrating robust generalization across diverse model families. In subsequent analyses, VisJudge refers to this best-performing variant unless otherwise specified. For analysis of how training data scale affects these results, see Appendix E.1.